OpenStudy (anonymous):

Tyler wants a pair of shoes that sells for $60.00. He uses two coupons when he buys these shoes. One coupon is for a 30% discount and the other coupon is for $10.00 off. The cashier takes the 30% discount first, and then subtracts $10.00. How much does Tyler pay for the pair of shoes? B. Would the price be different if the cashier has subtracted the $10.00 and the taken the 30% discount?

OpenStudy (anonymous):

pardon me if i m wrong i think you will first calculate 30% on $60 i.e 60*30/100 then subtract that amount from $60 and then subtract $10 from the remaining amount Tyler will pay $32 solution 60*30%=18 60-18=42 42-10=$32 is it ok...?

OpenStudy (kymber):